Minutes of the November 13, 2003 Tracker Meeting



      - Review old action items

      - Status of vibe test preparations

      - Status of T/V test preparations; schedule of TRR

      - Status of mid tray assembly preparations

      - Status of bottom tray assembly preparations; schedule

      - Status of MCM preproduction and GTRC V7

      - A.O.B.


New Action Items:

  1. Hytec still needs to send to Bari the vibration predictions in Excel form.  (closed)
  2. Tom: send torz+ M4 fasteners to Hytec.
  3. Tom: get the insert pull coupon data from COI and send to Hytec.
  4. Nanda: write up a schedule and plan for getting from the vibration test to the thermal-vacuum test.
  5. Jack, Nicola, Tom: resolve what thermocouples and wire to use.
  6. Erik: send to Plyform the step files for the mini-sidewall drawings.
  7. Nanda: find the status of coupon testing at Plyform and send to Tom and Robert.
  8. David Rich: clarify with Darren mass requirements on the tray panel drawings.
  9. Jim Martin: investigate availability of manpower for completing the flex-circuit cable design.


Old Outstanding Action Items (Sandro was not present so some of these did not get addressed in this meeting):

  1. Tom: discuss the new vibration test information from Italy with NASA.  (closed, Farhad agrees with what we are planning)
  2. Robert: request confirmation from Gunther about EGSE hardware purchases.  (closed)
  3. Sandro: send document drafts 1601 and 1918, as well as the update to 1584.  (still missing 1918)
  4. Sandro: complete inspection of the heavy cores.  (open)
  5. Jeff: send the airbill numbers to Sandro for the vibration fixture shipments.  (closed; fixtures in Pisa)
  6. Erik: send the vibration predictions to Bari in Excel format.  (open)
  7. Sandro: send to Tom information on torque striping paint that they have available.  (open)
  8. Sandro: propose a detailed schedule for the T/V test.  (open)
  9. Tom, BJ: finalize and release drawings for the bottom-tray closeouts. Keep COI stopped until this is ready. (closed for titanium, closeouts in progress)
  10. Tom: estimate delivery times for the titanium parts and bottom-tray closeouts.  (open)
  11. Sandro: write down a plan for integrating the EM tower into the T/V chamber and test setup. (open)
  12. Bari: update the T/V test plan and procedure.  (closed)


Erik reported that he received the static-test sidewall panels Monday.  Strain gauges are now bonded to them.  Starting this morning they are setting up for the first axis (z).  They will start pulling probably Friday morning.  Tom needs to send torx+ M4 fasteners to Erik.  Meanwhile Hytec has to use the offset cruciform heads, which are a pain to get back out.  Erik needs the coupon data from COI.  Tom will find out after the meeting.


Tom sent T/V test curves to Nicola yesterday from Jeff Wang.  Later today he should get curves for time to plateau for power testing.  The first step will take several hours, the rest much quicker.  The plan now is to do 6 total thermal cycles, including the power thermal balance test.  Jeff predicts that the center of the top trays will take 10 hours to go from cold to hot or hot to cold.  There was a discussion about definition of reaching one end or the other.  The proposal is to eliminate and long soak at the extremes, and just hit within a few degrees of the maximum range at all parts of the tower and then turn around.  With this Tom thinks we can do 5 cycles in about 4 days, plus 2 days for thermal balance.


Nanda said that he needs information to correlate this plan to particular thermocouples to inform Alenia.  That is, the plan needs to state for specific thermocouples what they must read to define having reached a temperature extreme.  Tom is still waiting for predictions on stabilization at different power levels during the thermal balance test.


Sandro sent out initial drawings of the thermal blanket and support frame.


Nanda said that he will write out a schedule and plan for getting from the vibration test to the T/V test.


The draft plan and procedure for the T/V test are on the Bari web page for review.  Tom said that he will get a better copy of the curves to put into the plan, into the Word document.


There was a discussion on what thermocouples and wire to use.  There may be some already at SLAC that could be used.  This did not reach resolution and needs to be taken up offline between Jack, Nicola, and Tom.  There is a total of 54 thermocouples.


Nanda reported on the EM tower status.  The sidewalls came in last night.  The cable attach fixtures is being modified to fit better and align to the cables.  Then they will start changing the sidewalls.  This should be finished by mid day tomorrow.  There are some missing screws for the sidewalls of the inner shipping container.  Tom says that these screws were left at G&A.  The first of these units has American screws but the design has since been changed to metric.  The plan is to take the EM tower to Alenia on Monday.


Plyform has cut the sidewall coupons.  They cannot get access to the mini-sidewall drawings.  Erik can send the step files.  Nanda will find out the status of coupon tests and send to us.  They will do the tests in the order compression, shear, then tension. 


Mid tray assembly preparations:  The web page for the PRR has been updated with an action item list.  Many of the items require looking directly at Plyform documentation.  Nanda said that it is difficult for him to be at Plyform in the next weeks.  However he may be able to meet up with Darren or Jerry for half a day at Plyform.  One of the actions is about mass requirements on drawings.  What is really required there?  A not to exceed value?  A minimum value?  Dave Rich will clarify this with Darren.


Erik said that titanium part machining took 8 to 10 weeks last time, so this time is starting to get critical.  We need all to respond on the release of the parts drawings.  Separate assembly drawings are in progress.  The flexures will be assembled into the corner brackets before shipping to Italy.


MCM preproduction:  The first lot of 6 is progressing with component assembly.  Results can be seen in the Tracker presentation from last Wednesday’s management meeting.  The second batch of 6 is getting bonded and will ship to SLAC Friday.  SLAC will CMM inspect them Monday and ship them back to Teledyne. There is a QA survey tomorrow at Teledyne.


Flex-circuit cables: Hartmut reported that one of the main issues was to get a required length from mechanical systems.  The IDD will be signed off at the end of the month; then we can have the PRR.  We do have an agreed upon length, so the first drawings are being finalized.  The schedule depends on the workloads of BJ and Thien.  We will make mini-tower cables as soon as the drawings are ready, and we will use them to check fit on the mini-tower.  Jerry is talking with Parlex about production.  The plan has been changing.  Now we will do the bending cold just before installation, because the conclusion is that even a hot bend done initially will relax away and be pointless added cost.  BJ has to modify his tool for bending cactus arms for the new 8mm radius.  It can be used with connectors installed.  Jim will look into the questions of manpower availability for flex-circuit cable design.